# CRAFT Your Dataset: Task-Specific Synthetic Dataset Generation Through Corpus Retrieval and Augmentation

### v1

@@ 6 Conclusion
In this work, we introduced CRAFT (Corpus Retrieval and Augmentation for Fine-Tuning), a framework for generating task-specific synthetic datasets grounded in text corpora. CRAFT requires only a small set of human-curated few-shot examples to bootstrap the creation of large-scale training data by leveraging existing corpora and instruction-tuned language models. Our experiments across multiple tasks, including biology, medicine, and commonsense question-answering, as well as summarization, demonstrate that models fine-tuned on CRAFT-generated datasets can match or outperform strong baselines, including instruction-tuned models and those trained on human-curated datasets. Notably, CRAFT-based models showed better generalization capabilities on out-of-domain datasets compared to models trained on human-curated data, highlighting the robustness of our approach.
@@ Acknowledgements
IZ and DE have been supported by the European Union’s Horizon 2020 research and innovation program under grant agreement No. 101135671 (TrustLLM). AK and HS have been funded by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) project SCHU 2246/14-1. IZ was also supported by a G-Research Travel Grant to present this work at the ELLIS Doctoral Symposium 2024 in Paris.
@@ A.1 Few-Shot Design
The BioQA few-shot text samples were compiled from a diverse range of sources, including textbooks (Alberts, 2017; Malmquist and Prescott, 2022; Wilkin and Brainard, 2016; Rye et al., 2016), openly accessible materials, and Encyclopedia Britannica. For MedQA, we primarily drew upon openly accessible websites, such as the National Institutes of Health, National Health Service, Food and Drug Administration, Mayo and Cleveland Clinics, and other medicine-related websites, to generate our few-shot text samples. The CSQA few-shot text samples were sourced from a variety of online resources, including blogs, articles, and other websites tailored to the specific topic at hand. From sources that are not openly accessible through websites, continuous text snippets were directly extracted and used as texts, while all other material was shortened, rephrased, or restructured by the authors. This process ensures that articles which may have been crawled through C4 (Raffel et al., 2020) do not produce exact matches at retrieval time. Since none of these texts have direct question-answer pairs available, the question, answer options, and the answer were generated by the authors for each sample. Table 2, step 1, shows an example.
The recipe generation few-shots were taken from blogs and other openly accessible websites. Their design always includes a simple instruction or question to cook a meal, as well as bulleted or numbered lists of ingredients and steps to produce the meal. Sometimes, the recipes on websites were described only as continuous text. In such cases, the authors added the list of ingredients and steps to create the few-shots. Similar to the QA few-shots, it is desirable to increase the vocabulary diversity for retrieval. Therefore, we made sure to cover a wide range of recipes in the few-shots. To design summarization few-shots, we collected a wide variety of texts from websites, blogs, and magazines, as well as specialized sources such as GitHub issues. Here, the few-shots feature a text, an instruction to create a summary, as well as the summary output. On some websites, summarized versions of the main text are given, while in other cases continuous or bulleted summaries were created by the authors.
A.2 Embedding Database
To create the embeddings, SentenceTransformers (Reimers and Gurevych, 2019) is used, specifically a MiniLM (Wang et al., 2020) version (multi-qa-MiniLM-L6-cos-v1) optimized for cosine-similarity search between two document pairs. This model creates 384D embeddings, which are stored in an HDF5 database (The HDF Group, 2002), that allows for native storage and retrieval of array-like data.
A.3 Corpora
@@
The Colossal Clean Crawled Corpus (Raffel et al., 2020, C4) dataset consists of approximately 750GB of English data, pre-cleaned by the creators to exclude non-informative texts. We use a 305GB subset of the dataset, which excludes not-safe-for-work or offensive content.
The English Wikipedia corpus comprises a diverse and high-quality collection of textual information on a wide range of topics. We use the January 2nd, 2024 dump of English Wikipedia, which we pre-process using WikiExtractor (Attardi, 2015) to extract clean text documents.
The Stack Exchange corpus (Flax Sentence Embeddings Team, 2021) features a structured format with title, body, and best-voted answer collected from the Stack Exchange network. The dataset encompasses 173 distinct sub-communities, covering both technical and non-technical topics.
The WikiHow corpus (Koupaee and Wang, 2018) presents information in a step-by-step instructional layout, making it a valuable resource for tasks such as summarization or recipe generation. Each document consists of a title, clear instructions, and accompanying text.
@@
After filtering out documents with fewer than 200 characters or more than 25,000 characters, the resulting datasets contain approximately 362 million documents from C4, 10.5 million documents from English Wikipedia, 9.5 million documents from Stack Exchange, and 190,000 documents from WikiHow. The resulting 383 million documents take up approximately 247GB of storage when GZIP compressed (Deutsch, 1996) and represented as 16-bit NumPy arrays (Harris et al., 2020).
A.4 Document Retrieval
To retrieve relevant human-written samples, we employ a two-step process that approximates a global similarity search. Due to the large size of the embedding database (383 million documents), a direct global pair-wise comparison is impractical.
In the first step, we divide the embedding database sequentially into subsections of approximately 350,000 documents each. We calculate cosine similarity between each subsection and the few-shot samples, and store the top 5% most similar documents in memory. This reduces the number of documents considered from 383 million to approximately 19 million. In the second step, we recalculate cosine similarity and perform traditional top-k similarity retrieval on the filtered documents. This yields the indices of the top-k most similar documents to the few-shot examples.
Optimizing solely for top-k similarity between individual few-shot samples and the embedding database may lead to limited variation and high similarity among the retrieved documents. Conversely, optimizing for similarity by averaging the embedding representations of all few-shot samples may prioritize a single dominant topic, potentially biasing the selection. To balance these approaches, we adopt a mixed similarity retrieval strategy, combining two methods: (i) 50% of samples are retrieved based on individual top-k similarity to each few-shot sample, ensuring a minimum number of similar samples for each few-shot and mitigating topic dominance, and (ii) the remaining 50% of samples are retrieved based on fully averaged similarity, which aims to help uncover more latent topics.
A.5 Task Sample Synthesis
To facilitate the creation of high-quality synthetic task samples, we align the text generation process to the few-shot design through in-context learning (Brown et al., 2020). This approach helps reduce issues like hallucinations (Lei et al., 2023; Wang et al., 2023a) and formatting errors, although it does not entirely prevent them. Specifically, three few-shot examples from the human-curated few-shots are dynamically and randomly sampled for each input prompt and corpus sample, interleaving them with the instruction to augment the text. Since each few-shot example includes a text similar in length to corpus samples, the three-shot prompting technique can result in input prompts that often exceed 10,000 tokens, sometimes reaching 20,000 tokens. For optimal performance, it is recommended to use models offering long context lengths.
In our experiments, all task samples are created using Mistral 7B Instruct v0.2 (Jiang et al., 2023) along with the few-shot task template shown in Table 2, step 3. During the generation process, vLLM (Kwon et al., 2023) is used with temperature 0.7, top-k (Fan et al., 2018) and top-p (Holtzman et al., 2020) sampling at 40 and 0.9, respectively. The maximum generation length was determined based on preliminary experiments with the goal of accommodating full-length sample generation. For all QA datasets we limited samples to 256 tokens, whereas recipe generation and summarization were limited at 1280 and 1536 tokens, respectively.
To enable quality control checks such as checking whether the QA datasets have the desired number of answer options, or whether the samples have been fully generated, we instruct the model to produce all task samples as valid JSON objects with a set number of keys and proper naming. We perform checks on each JSON object to ensure that the questions, instructions, and generative task answers exceed a reasonable minimum length, as well as that the answer option generated for the QA tasks contain valid letters. If no valid JSON is found, or keys are missing, or any formatting check fails, we discard the sample. Furthermore, we filter task samples that are either too similar to the few-shots or to other generated task samples using fuzzy string matching (Ranjith et al., 2022) at a token set similarity ratio of 0.85. To account for the removal of samples during deduplication, we recommend retrieving approximately twice as many corpus samples as the number of desired final synthetic task samples. A detailed breakdown of the retrieval and deduplication quantities for each task is provided in Appendix B.
A.6 Training and Optimization
For all experiments, low-rank adaptation (Hu et al., 2021, LoRA) fine-tuning is performed with 16-bit BrainFloat (Abadi et al., 2016) as the computation data type. All experiments are implemented using PyTorch (Paszke et al., 2019) and the Hugging Face packages (Wolf et al., 2020). For optimization, the adaptive momentum optimizer with decoupled weight decay (Loshchilov and Hutter, 2019) of 5% and a learning rate of $1\times 10^{-4}$ is employed. A linear learning rate schedule is applied with a warm-up ratio of 10%. Fine-tuning is performed for 3 epochs across all tasks and dataset size variations. When fine-tuning on only the human-curated few-shots, a batch size of 2 is adopted. In all other scenarios, fine-tuning is performed with a batch size of 16 or the equivalent with gradient accumulation steps. Following Dettmers et al. (2024), LoRA adapters are used on every linear layer, specifically on the query-, key-, value-, output-, gate-, up-, and down-projection matrices. We use a rank and $\alpha$ parameter of 64, the adapter’s dropout rate is set to 0.1, and the bias terms of the update matrices are deactivated. In total, this adds 2.3%, or about 160 million parameters, of the model’s 7 billion base parameters as LoRA adapters. The rest of the model remains frozen during fine-tuning, and the updated parameters will be merged with the base model after fine-tuning.
Appendix B Filtering Statistics
@@
Appendix C Dataset Statistics

### v2

@@ 7 Conclusion
In this work, we introduced CRAFT (Corpus Retrieval and Augmentation for Fine-Tuning), a framework for generating task-specific synthetic datasets grounded in text corpora. CRAFT requires only a small set of human-curated few-shot examples to bootstrap the creation of large-scale training data by leveraging existing corpora and instruction-tuned language models. Our experiments across multiple tasks, including biology, medicine, and commonsense question-answering, as well as summarization, demonstrate that models fine-tuned on CRAFT-generated datasets can match or outperform strong baselines, including instruction-tuned models and those trained on human-curated datasets. Notably, CRAFT-based models showed better generalization capabilities on out-of-domain datasets compared to models trained on human-curated data and maintained robustness to variations in the quality of the initial few-shot examples. Furthermore, while some fully synthetic methods such as Self-Instruct produce competitive results, CRAFT outperforms these approaches overall, offering a more scalable and reliable dataset generation framework.
@@ Acknowledgements
We thank the action editor, Tao Ge, and the anonymous reviewers at TACL for their helpful comments during the review process. IZ and DE have been supported by the European Union’s Horizon 2020 research and innovation program under grant agreement No. 101135671 (TrustLLM). AK and HS have been funded by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) project SCHU 2246/14-1. IZ acknowledges the EuroHPC Joint Undertaking for awarding access to MareNostrum5, hosted at Barcelona Supercomputing Center (BSC), Spain, under proposal No. EHPC-DEV-2024D12-031. IZ was also supported by a G-Research Travel Grant to present this work at the ELLIS Doctoral Symposium 2024 in Paris.
@@ A.1 Few-Shot Design
BioQA few-shot texts were drawn from diverse sources, including textbooks (Alberts, 2017; Malmquist and Prescott, 2022; Wilkin and Brainard, 2016; Rye et al., 2016), Encyclopedia Britannica, and openly accessible materials. MedQA examples were based on public content from health-related websites (e.g., NIH, NHS, FDA, Mayo Clinic, Cleveland Clinic). CSQA few-shots were compiled from blogs, articles, and other topical online sources. Recipe generation few-shots were sourced from blogs and public recipe sites, typically structured as an instruction or question with lists of ingredients and steps. If recipes were in continuous prose, the authors added structured elements manually. To ensure diverse retrieval, a wide range of dishes and vocabulary was included. Summarization few-shots were built from texts across websites, blogs, magazines, and GitHub issues. Each example included a full input text, a summarization instruction, and a corresponding summary. Where summaries were not available, they were authored to match realistic format and content.
A.2 Corpora
@@
C4 (Raffel et al., 2020) provides a broad web-crawled dataset. We use a 305GB filtered subset of the original 750GB corpus, excluding not-safe-for-work content. Wikipedia offers high-quality encyclopedic content. We use the English Wikipedia dump from January 2, 2024, processed with WikiExtractor (Attardi, 2015) to extract clean text. Stack Exchange (Flax Sentence Embeddings Team, 2021) includes QA-format documents across 173 communities, combining title, body, and top-voted answer. It spans technical and non-technical domains. WikiHow (Koupaee and Wang, 2018) features instructional content in a step-by-step format, useful for generative tasks such as recipe generation or summarization.
@@
After filtering documents to lengths between 200 and 25,000 characters, we retain 362M documents from C4, 10.5M from Wikipedia, 9.5M from Stack Exchange, and 190K from WikiHow. Combined, these 383M documents occupy 247GB when GZIP-compressed (Deutsch, 1996) and stored as 16-bit NumPy arrays (Harris et al., 2020).
A.3 Document Retrieval
We use a two-step retrieval strategy to approximate global similarity search over 383M embedded documents, avoiding expensive full pairwise comparisons.
First, the embedding database is split into sequential shards of 350K documents. For each shard, cosine similarity to the few-shot samples is computed, and the top 5% most similar documents are retained, reducing the candidate pool to 19M documents. Secondly, we perform a second round of cosine similarity and standard top-$k$ retrieval on this reduced set to obtain the final document matches.
To avoid redundancy or topical overfitting in retrieval, we combine two strategies: (i) 50% of samples are retrieved via top-$k$ similarity to each few-shot individually, and (ii) 50% via similarity to the averaged embedding of all few-shots to capture latent task structure.
A.4 Task Sample Synthesis
Synthetic task samples are generated via in-context learning (Brown et al., 2020), using three randomly sampled few-shot examples per prompt. These are interleaved with the instruction and a retrieved corpus document, following the template shown in Table 2, step 3. Because each few-shot includes a long text, input prompts often exceed 10,000 tokens (up to 20,000), requiring long-context models.
We use Mistral 7B Instruct v0.2 (Jiang et al., 2023) with vLLM (Kwon et al., 2023) for generation. Sampling is performed with temperature=0.7, top-kk=40 (Fan et al., 2018), and top-pp=0.9 (Holtzman et al., 2020). Maximum output lengths are capped at 256 tokens for QA, 1280 for recipes, and 1536 for summarization, based on empirical tuning.
To enable quality control, all outputs are formatted as JSON with fixed keys. We discard any samples with malformed structure, missing fields, or insufficient length. For QA tasks, we validate that answer options are complete and contain a valid label. To reduce redundancy, we filter out samples with high similarity to few-shots or other outputs using fuzzy string matching (Ranjith et al., 2022) with a token set ratio >0.85>0.85. We recommend retrieving roughly twice the desired number of corpus samples to account for filtering. Task-specific filtering statistics are detailed in Appendix D.
Appendix B Additional Results: Varying the Number of Few-Shots
Figure 7: Performance between CRAFT when initiated with 8 and 32 human-curated few-shot examples.
Appendix C Full-Length CRAFT-Generated Examples
C.1 Recipe Generation
Instruction: Prepare apple strudel using puff pastry and apples with sugar and raisins.
Ingredients: 5 medium apples, peeled, cored, and sliced; 1/2 cup granulated sugar; 1/2 cup raisins; 1 pound puff pastry, thawed; 1 egg, beaten; 1 tbsp water; powdered sugar, for dusting.
Steps: Cook the apples: In a saucepan, cook the sliced apples with sugar over medium heat until softened and caramelized. Combine apples and raisins: Transfer the cooked apples to a bowl and mix in the raisins. Prepare the pastry: Unroll the puff pastry on a lightly floured surface and place it on a baking sheet. Add apple filling: Place the apple filling in a line down the middle of the pastry, leaving a 2-inch border on all sides. Brush the edges: Brush the edges of the pastry with the beaten egg and water mixture. Fold the pastry: Fold the edges of the pastry over the apple filling, sealing it tightly. Bake the strudel: Brush the top with the egg mixture and bake at 375°F for 30–35 minutes, until golden brown. Cool and serve: Let cool slightly before dusting with powdered sugar and serving.
Instruction: Prepare a dish using items from your well-stocked pantry, freezer, and fridge.
Ingredients: Honey, Chipotle peppers in adobo sauce, Assorted olives, jalapenos, pepperoncinis etc., Flour (all-purpose, whole wheat, self-rising), Sugar, brown sugar, Powdered sugar, Baking powder, baking soda, extracts, yeast, Masa, Potatoes, onions, garlic, Dried beans, Stocks and broths (chicken, beef, vegetable), Jarred pesto and specialty relishes, chutneys, etc., Panko breadcrumbs, Kosher salt, black pepper, Worcestershire, Tabasco, olive oil, different vinegars, soy sauce, Longer lasting cheeses (cheddar, Parmesan, feta), Bacon, Corn tortillas, Beef, sausage (breakfast sausage, Italian sausage, chorizo), Dinner rolls, Pie crust, Frozen fruits (peaches, berries, cherries, etc.), Frozen vegetables (green beans, peas, lima beans, carrots, corn), Ready-made recipes, Chili, Homemade pumpkin puree
Steps: Choose a recipe or decide on a dish to make using the available ingredients. Prepare and gather ingredients from the pantry, freezer, and fridge. Cook the dish according to the recipe or your preferred method. Serve and enjoy!
Instruction: Create a salad or bowl by combining greens, proteins, vegetables, cheeses, fruits, and nuts with the specified dressing.
Ingredients: Chopped cucumber, Tomato, Bell pepper, Red onion, Feta cheese, Kalamata olives, Chopped green kale, Shredded chicken, Black beans, Crumbled cojita, Diced avocado, Spicy crispy pepitas, Lemon-cumin vinaigrette, Chopped Napa cabbage, Rice noodles, Julienned red bell peppers, Scallions, Shredded carrots, Cilantro, Peanuts, Zingy miso ginger vinaigrette, Crispy noodles, Sweet corn, Diced bell pepper, Tomato, Avocado, Red onion, Cilantro, Toasted pine nuts, Blue cheese vinaigrette, Dried cherries, Mild feta cheese, Swiss chard, Peppery baby arugula, Toasted pecans, Champagne vinaigrette, Fresh baby lettuce, Sweet green peas, Crunchy bacon shards, Savory white mushroom sauté, Asiago cheese, Tart vinaigrette, Grilled Bosc pears, Bacon crumbles, Toasted pecan halves, Tangy Maytag blue cheese, Tender greens, Ruby port vinaigrette, Frisée greens, Poached egg, Bacon lardons, Warm shallot vinaigrette, Baby spinach, Sautéed mushrooms, Bacon, Hardboiled egg pieces, Slivered red onion, Honey balsamic vinaigrette, Red cabbage, Romaine, Tomato, Corn, Green and red onion, Seasoned beef and chorizo, Cheddar and jack cheese, Cumin honey-lime vinaigrette, Duck confit, Hoisin and hot sauce, Peanuts, Cabbage, Scallions, Carrots, Crisp jicama shells, Assorted fruits, Scoop of ice cream
Steps: Chop or prepare all vegetables, fruits, and proteins as specified in the text. Combine the greens in a large bowl or on individual plates. Add proteins, vegetables, fruits, cheeses, and nuts to the greens. Drizzle with the specified dressing or vinaigrette. Toss gently to evenly coat the salad or bowl. Serve and enjoy!
C.2 Summarization
Instruction: Summarize the text below, focusing on Bill Ackman’s investment in Procter & Gamble and his intentions for the company.
Input Text: Bill Ackman, a well-known hedge fund manager, believes that Procter & Gamble (P&G) could be worth more than its current share price due to poor marketing and pricing strategies. Ackman holds around 1% of the company’s shares and has been actively advocating for changes. During a presentation, he mentioned that he spends 25% of his time on his investment in P&G. Ackman has been critical of the present management and its financial performance since disclosing his stake last year. It remains to be seen whether the company will welcome Ackman’s involvement more enthusiastically than in the past. In the past, P&G has been targeted by activists after an event-driven turn in the share price. Ackman has not specified who he would like to see as the new CEO but has mentioned that potential changes could lead to savings, marketing efficiencies, and operating leverage. He did not mention any particular candidates during his presentation.
Generated Summary: Bill Ackman, a hedge fund manager, believes that Procter & Gamble’s (P&G) stock could be worth more than its current price due to poor marketing and pricing strategies. He holds around 1% of the company’s shares and is actively pushing for changes. Ackman has been vocal about his dissatisfaction with the current management and has spent significant time advocating for improvements. He has not specified who he would like to see as the new CEO but has mentioned that potential changes could lead to savings, marketing efficiencies, and operating leverage.
Appendix D Filtering Statistics
@@
Appendix E Dataset Statistics